# Are Vision Language Models Texture or Shape Biased and Can We Steer Them?

### v1

@@ 6 Conclusion
In this study, we investigated texture/shape bias in VLMs. Through the lens of this specific bias, we are able to assess whether visual biases are inherited from vision encoders or modified through text processing. We find that the latter is the case: VLMs are often more shape-biased than their vision-only backbones. Beyond that, an intriguing aspect of VLMs is that the task/response can be steered through not just vision but also through language. Our experiments have demonstrated that VLMs have learned a multimodal association between the terms “shape”/“texture”, their synonyms, and their respective visual concepts to some extent. Taken together, we see that when it comes to multimodal models, the sum is greater than the parts, as a visual bias (texture/shape bias) can be steered to a certain degree through both manual and automated prompt engineering, alone.
Similar to the active research line on prompt engineering for CLIP [75, 76, 77, 78], it remains to be seen if prompts exist that can steer texture/shape bias to the same extent as vision can. While this aligns with human behavior in psychophysical studies [1], it remains unclear if this is a reflection of a learned bias from human-annotated data or a spurious correlation. We plan to explore this further in future studies.
Lastly, despite experiments on an extensive and diverse collection of VLMs, we have seen a (surprisingly) homogenous landscape in terms of texture/shape bias, steerability, and error consistency. We are curious to see how our insights apply under more radical training/architecture changes and to other visual biases. We encourage the community to provide more open-source implementations without which studies like ours would not be feasible.
Limitations.
Our study, like most studies, is not free of limitations. Although we utilized a diverse array of VLMs, there is a possibility that we overlooked significant outliers. Moreover, the applicability of our findings to future VLMs remains uncertain; however, we believe our results effectively reflect the current landscape. A potential bias introduced in our study stems from the prompts we employed. Different prompts could have yielded disparate results, and certain models might have performed differently, particularly in Visual Question Answering (VQA) tasks. While we mitigated this by utilizing simple, widely used prompts (refer to the appendix for an exploration of alternatives), the possibility of better choices remains open for future investigation.
@@ Appendix A Overview of VLMs
Qwen-VL-Chat [63]
Adds vision capabilities to Qwen-7B [79]. We set a repetition penalty of $1.2$ for this model.
Qwen-VL Plus/Max [26]
AliBabas’s proprietary larger variants of Qwen-VL-Chat. Access only via API.
CogAgent [64]
A special model for interaction with graphical user interfaces (GUIs) at high-resolution.
CogVLM [65]
Adds "trainable visual expert module" in LLM layers to combine vision and language.
Emu2 [66]
The 37B model claims "strong multimodal in-context learning abilities".
InstructBLIP [23]
Connects frozen vision encoders and LLMs through a trainable Q-Former. Uses Vicuna or FLAN-T5 as LLMs.
LLaVA v1.5 [61]
Improvements of LLaVA with modifications on the image encoder, the projector, and task-specific data. Uses Vicuna-7/13B as LLM.
LLaVA-NeXT [62]
Successor of LLaVA v1.5 supporting higher resolutions through patching, and using better SFT training data for training, claiming “improved reasoning, OCR, and world knowledge” [62]. The 34B version switches from Vicuna-7/13B to Nous Hermes 2 Yi 34B.
MoE-LLaVA v1.5 [67]
Variants of LLaVA v1.5 employing 4 sparsely activated Mixture-of-Experts (MoE), and smaller LLMs (Qwen, Phi-2, StableLM).
LLaVA-RLHF [69]
Variants of LLaVA v1.5 aligned with Factually Augmented RLHF (Fact-RLHF) [69].
UForm-Gen Chat [68]
A small (1.5B) model for VQA and image captioning finetuned for multimodal chat.
Gemini 1.0 Pro Vision [25]
Google’s proprietary multi-modal model based on the Gemini Pro LLM. Access only via API.
InternVL Chat 1.1/1.2+ [2]
An open-source effort to provide an alternative to ViT-22B [80]. V1.1 is based on a 6B ViT and Vicuna-13B, V1.2+ uses Nous Hermes 2 Yi 34B as LLM including additional SFT on 10x more data.
GPT-4V (Preview) [12]
OpenAI’s proprietary multi-modal model based on the GPT-4 LLM. Access only via API. Often considered to be the most powerful model.
For our main analysis, we prompt all models at the default generation parameters (e.g., temperature) unless stated otherwise. Appendix F shows why this is not an issue.
@@ Appendix B Detailed Results Table
Tab. 2 shows the shape bias and accuracy for the VQA and Image Captioning task (see Fig. 2 in the main paper for a visualization). For the open-ended Image Captioning responses, we additionally provide evaluations through an LLM (see Sec. 3). These include the number of generated tokens (to measure how effective our “Keep […] short.” instruction is), the ratio of responses where exactly one class was detected (single class ratio), and the ratio of responses that do not refer to any description (generic ratio). In Appendix I/Tab. 9, we provide ablations on Image Captioning under the removal of generic responses.
@@
Gemini 1.0 Pro Vision [25] 64.1 82.33 63.2 68.00 18.9 63.0 32.3
GPT-4V (Preview) [12] 47.9 69.75 53.6 52.67 44.8 37.2 60.4
Qwen-VL Plus [26] 64.8 82.92 67.9 65.50 21.9 59.2 36.0
Qwen-VL Max [26] 62.4 85.50 69.7 68.50 151.9 52.1 41.0
Qwen-VL Chat [63] - - 38.2 67.42 27.3 59.1 33.2
InternVL Chat 1.1 [2] 68.3 89.33 73.2 75.58 16.9 74.9 19.4
InternVL Chat 1.2+ [2] 61.1 90.83 61.3 82.42 15.8 80.4 11.4
LLaVA v1.5 7B [61] 61.4 80.75 61.4 76.08 12.1 73.8 19.2
LLaVA v1.5 13B [61] 64.1 80.25 62.7 75.58 28.9 65.8 23.8
LLaVA-RLHF 7B [69] 61.7 68.08 63.0 71.83 47.9 65.1 24.7
LLaVA-RLHF 13B [69] 63.4 80.42 62.3 73.25 38.3 64.7 27.7
LLaVA-NeXT 7B [62] 59.2 82.58 64.0 65.08 20.2 55.5 39.5
LLaVA-NeXT 13B [62] 57.2 83.42 63.5 65.25 48.8 52.6 40.9
LLaVA-NeXT 34B [62] 56.0 73.83 66.2 57.50 93.4 36.2 59.1
MoE-LLaVA-StableLM [67] 59.1 80.08 63.0 73.92 24.1 67.4 21.6
MoE-LLaVA-Qwen [67] 62.9 59.50 63.2 75.33 13.3 69.4 20.7
MoE-LLaVA-Phi2 [67] 59.6 82.33 61.1 75.42 34.9 67.0 18.6
InstructBLIP Flan-T5-xl [23] 68.2 79.58 67.1 81.50 116.7 57.0 22.3
InstructBLIP Vicuna-7B [23] 73.8 72.25 67.7 80.67 94.0 60.9 28.0
Emu2-Chat [66] 52.9 75.08 59.6 65.00 13.6 63.0 34.0
CogAgent Chat [64] - - 67.4 60.33 40.1 49.6 47.7
CogVLM Chat [65] - - 57.6 66.58 35.8 53.2 40.1
UForm Gen Chat [68] - - 38.8 64.50 30.2 59.3 33.0
Appendix C Prompting
@@ C.2 Exploration of Alternative Prompts
In initial testing, we found that the choice of prompts affects the eventual results and has the potential to inevitably influence our study. Thus, in an effort to address this, we extensively evaluated our models with multiple different prompting techniques, used in literature [61, 23] and chose the best one. Our prompt for VQA is inspired by LLaVA’s prompts for multiple-choice questions  [^1] https://github.com/haotian-liu/LLaVA/blob/main/docs/Evaluation.md [Online; accessed 6. Mar. 2024]. In an additional experiment (Tab. 3), we ablated alternative prompts on LLaVA-NeXT 7B [62]. We change or use an empty VQA_INSTRUCTION and change options to CLIP-style options ("X. a photo of a {class}"). However, we only observed a minor fluctuation in accuracy and shape bias and no significant effects. Our default prompt delivers the best accuracy and is, thus, our preferred choice.
@@
"Which option best describes the image? [...]" (default)
59.2 82.58
Default with CLIP-style options
59.5 81.92
@@
with CLIP-style options
59.4 80.17
Empty instruction (just options)
59.5 81.33
Our image captioning prompt is a reformulation of the VQA prompt ("Which option best describes the image?" $\rightarrow$ "Describe the image."). In the following, we ablate if the suffix ("Keep your response short.") may have interfered with our results. Additionally, we tested an alternative suffix that explicitly asks for more details on LLaVA-NeXT 7B [62]. The results for the former investigation in Tab. 4, show that our suffix indeed did not heavily bias the results in terms of shape bias. While adding the suffix leads to an impact in accuracy, it reduces the ratio of generic descriptions (not referring to any class) and has on average almost 4x fewer tokens resulting in significantly faster inference. Switching the suffix to "Be precise." increases shape bias, but at the same time also increases the number of generated tokens and worryingly the ratio of generic responses. Overall, we find that captioning prompts are more fragile, but our chosen default prompt provides an intriguing balance. For all ablated prompts, we find that the shape bias is higher than in VQA.
@@
"Describe the image. Keep your response short." (default)
$64.0$ $65.08$ $55.5$ $39.5$
@@ C.3 Biased Prompts
We have retrieved the following strong synonyms from Thesaurus [73, 74].
@@
Appendix D LLM-based Response Extraction
We rely on an LLM to extract labels from the generated captions for the Image Captioning task. For every generated description we instruct Nous-Hermes-2-
Mixtral-8x7B-DPO
[54] with the following prompt:
@@ Appendix D LLM-based Response Extraction
Appendix E Automated Prompt Search
Loosely inspired by [45], we utilized an LLM to optimize prompts. we switched to Mixtral-8x7B-Instruct-v0.1222https://huggingface.co/mistralai/Mixtral-8x7B-Instruct-v0.1 [Online; accessed 6. Mar. 2024], as it performed better than the Nous Hermes version in early tests. The results shown in Sec. 5.2/Fig. 6 are a summary of multiple prompts that we tried in numerous multi-round conversations. We instruct the model to provide a prompt in a new line starting with “PROMPT: ” that we then extract and automatically evaluate. Afterwards, we return the results to the LLM and ask it to generate the next prompt.
@@ Appendix E Automated Prompt Search
Table 5: Example conversation for LLM-guided Prompt Search. $*{}^{*}$ock conversation.
Your task is to help me design an ideal image classification prompt for a vision/language model to MINIMIZE the shape bias without significantly hurting accuracy. The provided photos show natural objects modified to contain shape and texture of conflicting object classes. E.g., it could be a photo of an elephant (shape) with dog fur (texture). The model should classify the image as ’dog’ based on the texture of the object and ignore the shape. The model’s accuracy is the percentage of correctly classified images. The shape bias is the ratio of how often the model classified based on shape over texture. You can test your prompt by outputting a single new line starting with ’PROMPT: ’. Do not list options - the system will provide them automatically. Try to keep the prompt as short and simple as possible but be creative. It might be reasonable to summarize insights of previous attempts and to outline your goals before responding with a new prompt, but make sure that only the prompt starts with ’PROMPT:’. In response to the prompt you will be told the accuracy and shape bias. Then you will refine the prompt and we will continue until I say stop. Let’s go!
PROMPT: Which option best describes the image?
Prompt: […], Accuracy: 82.58 %, Shape Bias: 59.43 %. What is your next prompt?
Assistant
To minimize the shape bias while maintaining accuracy, let’s encourage the model to focus more on textures than shapes. Here’s my next prompt:
PROMPT: Based on the surface characteristics and patterns, which category does the image most resemble?
User
Prompt: […], Accuracy: 78.25 %, Shape Bias: 50.9 %. Can you improve this?
$\cdots$
Assistant
I have provided various prompts and explored different approaches to minimize the shape bias while maintaining high accuracy. The current optimal prompt is: […] Here are some additional suggestions for minimizing the shape bias: […]
These methods may require more computational resources and time investment. If you have other questions or need guidance on different topics, please don’t hesitate to reach out. Wishing you the best in your continued efforts!
At this point, I believe that we have explored various approaches, and the current prompt is the best-performing one among those tested. If you decide to pursue any of the mentioned alternatives, I am confident that you will find ways to improve the model’s performance and minimize the shape bias. Good luck with your continued work!
As our conversation has concluded, I would like to thank you for engaging in this exercise, and I hope I have been helpful in providing prompts and suggestions. If you have other questions or need guidance on different topics in the future, please don’t hesitate to reach out. Wishing you the best in your projects and research!
@@
Appendix F Ablation of Temperature Scaling
We are interested in determining if generation parameters can influence the behavior of shape bias. Generally, VLMs only expose a few controllable parameters but all offer some form of stochastic sampling of tokens, often via temperature scaling of the token logits. Most models default to low-temperature settings (or settle for a greedy token strategy) which is more correlated with precise answers and reasonable for VQA. On the contrary, higher temperatures are correlated with more creative outputs and eventually token gibberish at extreme values. In general, temperature scaling also results in better-calibrated models [81].

### v2

@@ 6 Conclusion
In our study, we investigated the processing of visual cues in multi-modal models, specifically if language can be used to change the visual perception of these models.
We acknowledge that the broader research question extends beyond the scope of any single study. Thus, our work focuses on a specific, well-defined example of a visual bias: the texture/shape bias (Geirhos et al., 2019) and investigates how language can be used to favor texture or shape in their predictions. Furthermore, we extend our analysis beyond texture/shape bias to include evaluations of low/high-frequency bias, demonstrating that the steerability of a model’s bias is not limited to a single characteristic.
Indeed, we were able to show that visual cues are influenced by language. Utilizing this finding, we can show an intriguing aspect of VLMs: biases in the response can be steered through simple natural language prompts offering a form of alignment. This unique trait differentiates LLM-based from task-specific and uni-modal models. While this form of steering was not able to fundamentally change the utilized cue for the evaluated biases [^3] Interestingly, our observed behavior of limited steerability has similarities to human perception experiments conducted by Geirhos et al. (2019). In their control experiments, humans were either instructed to identify the shape while ignoring the texture or conversely to identify the texture while ignoring the shape. This ”human prompt steering” worked, but only to a certain extent: When humans were tasked to ignore the shape, the human shape bias decreased from $96\%$ (neutral instruction) only to approx. $70\%$ shape bias (texture-biased instruction). Our tested VLMs behave somewhat similarly: their visual shape bias can be steered through prompting, but it appears hard for them to completely go against their default visual bias., it comes at almost no impact in accuracy, and most importantly does not require any retraining of the model. In fact, many attempts at steering shape bias in training have yielded worse results through more expensive methods Li et al. (2021); Lukasik et al. (2023). Instead, we provide a simple and intuitive way for practitioners to adjust the output beyond the inductive bias at runtime with minimal effort.
Taken together, we can indeed talk VLMs into seeing the world differently.
Limitations. Even though we utilized a diverse array of VLMs, there is a possibility that different models would lead to different conclusions. We believe our results provide a fair reflection of the current VLM landscape, but radically different VLM architectures may lead to changes in bias mechanics and consequently alter our findings. Further, not all models have been behaving as expected in our study: Given that GPT-4V often achieves SOTA performance and is considered an important baseline, it has surprisingly poor accuracy in both VQA and image captioning tasks compared to most other models - mostly due to refusal to answer which affected 131/1280 VQA conversations, i.e., roughly 10%. This is substantially higher than the refusal rate of all other models ($<1\%$). It is worth noting that refusal rates do not affect the shape bias measurement. GPT-4V is also the model with the largest amount of generic image captions ($60.4\%$). Additionally, we acknowledge that other prompts may have led to better results, however, the result is noteworthy, as the other VLMs mostly behave well under the same prompts. Overall, prompting is a potential source of bias in our study. Different prompts could have yielded different results, and certain models might have performed differently, particularly in Visual Question Answering (VQA) tasks. While we mitigated this by utilizing simple, widely used prompts, other choices remain to be explored in future investigations. We provide a brief exploration of alternatives in Sec. C.2.
Acknowledgments
PG and JK acknowledge financial support by the German Federal Ministry of Education and Research (BMBF) in the program “Forschung an Fachhochschulen in Kooperation mit Unternehmen (FH-Kooperativ)” within the joint project LLMpraxis under grant 13FH622KX2. PG is additionally supported by the German Federal Ministry of Education and Research (BMBF) project STCL - 01IS22067. JL acknowledges support from the Lamarr Institute for Machine Learning and Artificial Intelligence. JL, SJ, and MK acknowledge support by the German Research Foundation research unit 5336 “Learning to Sense”.
We thank Priyank Jaini for his valuable feedback.
Reproducibility Statement
We used open-source data and models, with detailed descriptions of our evaluation process provided in the appendix. The source code, evaluation results containing model answers for each sample, prompts generated by the automated prompt search, and the frequency-cue-conflict dataset, along with its creation scripts, are available at: https://github.com/paulgavrikov/vlm_shapebias.
@@ Appendix A Overview of VLMs
Qwen-VL-Chat Bai et al. (2023b) Adds vision capabilities to Qwen-7B Bai et al. (2023a). We set a repetition penalty of $1.2$ for this model.
Qwen-VL Plus/Max Qwen Team (2024) AliBaba’s proprietary larger variants of Qwen-VL-Chat. Access only via API.
CogAgent Hong et al. (2023) A special model for interaction with graphical user interfaces (GUIs) at high-resolution.
CogVLM Wang et al. (2023) Adds "trainable visual expert module" in LLM layers to combine vision and language.
Emu2 Sun et al. (2023a) The 37B model claims "strong multi-modal in-context learning abilities".
InstructBLIP Dai et al. (2023) Connects frozen vision encoders and LLMs through a trainable Q-Former. Uses Vicuna or FLAN-T5 as LLMs.
LLaVA v1.5 Liu et al. (2023a) Improvements of LLaVA with modifications on the image encoder, the projector, and task-specific data. Uses Vicuna-7/13B as LLM.
LLaVA-NeXT Liu et al. (2024) Successor of LLaVA v1.5 supporting higher resolutions through patching, and using better SFT training data for training, claiming “improved reasoning, OCR, and world knowledge” Liu et al. (2024). The 34B version switches from Vicuna-7/13B to Nous Hermes 2 Yi 34B.
MoE-LLaVA v1.5 Lin et al. (2024) Variants of LLaVA v1.5 employing 4 sparsely activated Mixture-of-Experts (MoE), and smaller LLMs (Qwen, Phi-2, StableLM).
LLaVA-RLHF Sun et al. (2023c) Variants of LLaVA v1.5 aligned with Factually Augmented RLHF (Fact-RLHF) Sun et al. (2023c).
UForm-Gen Chat Kim et al. (2023) A small (1.5B) model for VQA and image captioning finetuned for multi-modal chat.
Gemini 1.0 Pro Vision Gemini Team (2023) Google’s proprietary multi-modal model based on the Gemini Pro LLM. Access only via API.
InternVL Chat 1.1/1.2+ Chen et al. (2024) An open-source effort to provide an alternative to ViT-22B Dehghani et al. (2023). V1.1 is based on a 6B ViT and Vicuna-13B, V1.2+ uses Nous Hermes 2 Yi 34B as LLM including additional SFT on 10x more data.
GPT-4V (Preview) OpenAI (2023) OpenAI’s proprietary multi-modal model based on the GPT-4 LLM. Access only via API. Often considered to be the most powerful model.
For our main analysis, we prompt all models at the default generation parameters (e.g., temperature) unless stated otherwise. Appendix H shows why this is not an issue.
@@ Appendix B Detailed Results Table
Tab. 4 shows the shape bias and accuracy for the VQA and Image Captioning task (see Fig. 2 in the main paper for a visualization). For the open-ended Image Captioning responses, we additionally provide evaluations through an LLM (see Sec. 3). These include the number of generated tokens (to measure how effective our “Keep […] short.” instruction is), the ratio of responses where exactly one class was detected (single class ratio), and the ratio of responses that do not refer to any description (generic ratio). In Appendix L/Tab. 9, we provide ablations on Image Captioning under the removal of generic responses.
Further, we make the following observations:
Which models are the most shape-biased?
The strongest shape bias is observed in InstructBLIP Vicuna-7B Dai et al. (2023) for VQA, but the model generally shows a lower accuracy compared to other models. A more accurate model is InternVL-Chat 1.1 Chen et al. (2024) which ranks second place for VQA but first for captioning.
Does LLM scale matter?
LLM capacity does not seem to correlate with shape bias and unpredictably skews the shape bias by a few percent in each way as can be seen in Qwen-VL, LLaVA v1.5/NeXT/RLHF, or InternVL. Similarly, the overall largest models do not have the highest shape bias. However, following the overall general trend, in our experiments, we also found that scale usually improves accuracy.
Does RLHF align shape bias?
RLHF-tuned VLMs are still rare at this point and we only have three samples. On both LLaVA-RLHF Sun et al. (2023c) models we see no changes in comparison to the default LLaVA models. GPT-4V OpenAI (2023) (though it is unclear if vision was also RLHF trained) shows one of the lowest shape biases in our study, but we do not know how the base model ranks. Overall it is hard to derive a conclusive answer, but it seems that at least RLHF does not necessarily guarantee an alignment of visual preferences.
@@
Gemini 1.0 Pro Vision Gemini Team (2023) 64.1 82.33 63.2 68.00 18.9 63.0 32.3
GPT-4V (Preview) OpenAI (2023) 47.9 69.75 53.6 52.67 44.8 37.2 60.4
Qwen-VL Plus Qwen Team (2024) 64.8 82.92 67.9 65.50 21.9 59.2 36.0
Qwen-VL Max Qwen Team (2024) 62.4 85.50 69.7 68.50 151.9 52.1 41.0
Qwen-VL Chat Bai et al. (2023b) - - 38.2 67.42 27.3 59.1 33.2
InternVL Chat 1.1 Chen et al. (2024) 68.3 89.33 73.2 75.58 16.9 74.9 19.4
InternVL Chat 1.2+ Chen et al. (2024) 61.1 90.83 61.3 82.42 15.8 80.4 11.4
LLaVA v1.5 7B Liu et al. (2023a) 61.4 80.75 61.4 76.08 12.1 73.8 19.2
LLaVA v1.5 13B Liu et al. (2023a) 64.1 80.25 62.7 75.58 28.9 65.8 23.8
LLaVA-RLHF 7B Sun et al. (2023c) 61.7 68.08 63.0 71.83 47.9 65.1 24.7
LLaVA-RLHF 13B Sun et al. (2023c) 63.4 80.42 62.3 73.25 38.3 64.7 27.7
LLaVA-NeXT 7B Liu et al. (2024) 59.2 82.58 64.0 65.08 20.2 55.5 39.5
LLaVA-NeXT 13B Liu et al. (2024) 57.2 83.42 63.5 65.25 48.8 52.6 40.9
LLaVA-NeXT 34B Liu et al. (2024) 56.0 73.83 66.2 57.50 93.4 36.2 59.1
MoE-LLaVA-StableLM Lin et al. (2024) 59.1 80.08 63.0 73.92 24.1 67.4 21.6
MoE-LLaVA-Qwen Lin et al. (2024) 62.9 59.50 63.2 75.33 13.3 69.4 20.7
MoE-LLaVA-Phi2 Lin et al. (2024) 59.6 82.33 61.1 75.42 34.9 67.0 18.6
InstructBLIP Flan-T5-xl Dai et al. (2023) 68.2 79.58 67.1 81.50 116.7 57.0 22.3
InstructBLIP Vicuna-7B Dai et al. (2023) 73.8 72.25 67.7 80.67 94.0 60.9 28.0
Emu2-Chat Sun et al. (2023a) 52.9 75.08 59.6 65.00 13.6 63.0 34.0
CogAgent Chat Hong et al. (2023) - - 67.4 60.33 40.1 49.6 47.7
CogVLM Chat Wang et al. (2023) - - 57.6 66.58 35.8 53.2 40.1
UForm Gen Chat Kim et al. (2023) - - 38.8 64.50 30.2 59.3 33.0
Appendix C Overview of Prompts
@@ C.2 Exploration of Alternative Prompts
In initial testing, we found that the choice of prompts affects the eventual results and has the potential to inevitably influence our study. Thus, in an effort to address this, we extensively evaluated our models with multiple different prompting techniques, used in literature Liu et al. (2023a); Dai et al. (2023) and chose the best one. Our prompt for VQA is inspired by LLaVA’s prompts for multiple-choice questions  [^4] https://github.com/haotian-liu/LLaVA/blob/main/docs/Evaluation.md [Online; accessed 6. Mar. 2024]. In an additional experiment (Tab. 5), we ablated alternative prompts on LLaVA-NeXT 7B Liu et al. (2024). We change or use an empty VQA_INSTRUCTION and change options to CLIP-style options ("X. a photo of a {class}"). However, we only observed a minor fluctuation in accuracy and shape bias and no significant effects. Our default prompt delivers the best accuracy and is, thus, our preferred choice.
@@
"Which option best describes the image? [...]" (default) 59.2 82.58
Default with CLIP-style options 59.5 81.92
@@
with CLIP-style options 59.4 80.17
Empty instruction (just options) 59.5 81.33
Our image captioning prompt is a reformulation of the VQA prompt ("Which option best describes the image?" $\rightarrow$ "Describe the image."). In the following, we ablate if the suffix ("Keep your response short.") may have interfered with our results. Additionally, we tested an alternative suffix that explicitly asks for more details on LLaVA-NeXT 7B Liu et al. (2024). The results for the former investigation in Tab. 6, show that our suffix indeed did not heavily bias the results in terms of shape bias. While adding the suffix leads to an impact in accuracy, it reduces the ratio of generic descriptions (not referring to any class) and has on average almost 4x fewer tokens resulting in significantly faster inference. Switching the suffix to "Be precise." increases shape bias, but at the same time also increases the number of generated tokens and worryingly the ratio of generic responses. Overall, we find that captioning prompts are more fragile, but our chosen default prompt provides an intriguing balance. For all ablated prompts, we find that the shape bias is higher than in VQA.
@@
"Describe the image. Keep your response short." (default) $64.0$ $65.08$ $55.5$ $39.5$
@@ C.3 Biased Prompts
We have retrieved the following strong synonyms from Thesaurus Dictionary.com (2024b; a).
@@
Appendix D LLM-based Response Extraction for Captions
We rely on an LLM to extract labels from the generated captions for the Image Captioning task. For every generated description we instruct Nous-Hermes-2-Mixtral-8x7B-DPO Teknium et al. (2024) with the following prompt:
@@ Appendix D LLM-based Response Extraction for Captions
Appendix E Details on Automated Prompt Search
Loosely inspired by Yang et al. (2024), we utilized an LLM to optimize prompts. we switched to Mixtral-8x7B-Instruct-v0.1555https://huggingface.co/mistralai/Mixtral-8x7B-Instruct-v0.1 [Online; accessed 6. Mar. 2024], as it performed better than the Nous Hermes version in early tests. The results shown in Sec. 5.1/Fig. 4 are a summary of multiple prompts that we tried in numerous multi-round conversations. In Fig. 6 we show the impact on accuracy: in most cases the accuracy remains similar or even improves. Of course, outliers where accuracy severely decreases exist.
Figure 6: Shape bias vs. accuracy on cue-conflict in comparison for LLM generated prompts. Black points denote the default instruction.
We instruct the model to provide a prompt in a new line starting with “PROMPT: ” that we then extract and automatically evaluate. Afterwards, we return the results to the LLM and ask it to generate the next prompt.
@@ Appendix E Details on Automated Prompt Search
Appendix F Steering the Texture/Shape Bias in Vision
Earlier work demonstrated that ImageNet-models can still detect objects even if the image is split into patches and shuffled Zhang & Zhu (2019); Shi et al. (2020); Naseer et al. (2021). As patch size decreases, the operation is destroying more global shape information, yet retaining local texture information. We utilize this technique to significantly increase texture bias. Oppositely, to increase the shape bias we experiment with added Gaussian noise to inputs. This is loosely inspired by applying “diffusion-like noise” during training (and inference) which has been shown to drastically improve the shape bias of ImageNet-ResNets Jaini et al. (2024). However, we only apply the noise during inference and use a more simplistic approach by adding $\mathcal{N}(0,\sigma^{2})$ noise to all channels, consecutively clamping values to $[0,1]$. We visualize the effects on one cue-conflict sample in Fig. 7.
We show results on LLaVA-NeXT 7B in Fig. 8. Adding noise results increases the VLMs shape bias up to $89.5\%$ at $\sigma^{2}=0.5$, and patch shuffling decreases shape bias (increases texture bias) to $8.4\%$ at $28\times 28$ patches. In both cases, the bias is indeed steered up to a certain threshold: the accuracy on one cue (texture or shape accuracy) increases whereas the accuracy on the other decreases. Beyond a specific point, the operation is destroying one cue entirely and can no longer be considered steering. Further, this form of steering comes at a cost in accuracy - yet, all results are still well beyond random chance.
Figure 8: Image preprocessing can strongly steer texture/shape bias. Left: Shuffling image patches with decreasing patch size results in a strong texture bias. Right: Increasing Gaussian noise introduces a strong shape bias.
Inspired by these strong results, we repeat the experiments on the naturally more shape-biased and larger InternVL-Chat 1.1 Chen et al. (2024). In this model, we can further extend the range to $91.7\%$ shape bias ($\sigma^{2}=0.3$), and down to $6.1\%$ ($28\times 28$ patches).
Appendix G Frequency-cue-conflict
We create 1,200 samples belonging to 16 ImageNet-super-categories. The stimuli are generated by blending the low and high-frequency components of two differently labeled images. Specifically, we select two random ImageNet samples from the 16-class subset Geirhos et al. (2019) with conflicting labels. Each image is then converted to grayscale by selecting the L channel, resized to $256$ px on the shortest edge while preserving the aspect ratio, and finally center-cropped to $224\times 224$ px. We finally blend the two images by using 30% of the low-frequency components from one image and 70% of the high-frequency components from the other. The resulting stimuli are saved as JPEG with 100% quality. Fig. 9(b) shows some examples from the dataset.
Refer to caption Refer to caption Refer to caption Refer to caption
Shape: cat dog elephant truck
Texture: bottle clock dog chair
((a)) texture/shape-cue-conflict (Geirhos et al., 2019)
Refer to caption Refer to caption Refer to caption Refer to caption
LF: airplane bird clock elephant
HF: bottle boat bicycle truck
((b)) frequency-cue-conflict
Figure 9: Examples from the cue-conflict datasets. Images are constructed from two conflicting features cues: (a) texture/shape-cue-conflict (Geirhos et al., 2019) conflicting shape and texture; (b) frequency-cue-conflict conflicting high- (HF) and low-frequency (LF). Please zoom in for details.
Appendix H Ablation of Temperature Scaling
We are interested in determining if generation parameters can influence the behavior of shape bias. Generally, VLMs only expose a few controllable parameters but all offer some form of stochastic sampling of tokens, often via temperature scaling of the token logits. Most models default to low-temperature settings (or settle for a greedy token strategy) which is more correlated with precise answers and reasonable for VQA. On the contrary, higher temperatures are correlated with more creative outputs and eventually token gibberish at extreme values. In general, temperature scaling also results in better-calibrated models Guo et al. (2017).